OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x – 4| + 1 = 8 Step 2: |x – 4| = 8 – 1 Step 3: |x – 4| = 7 Which of the following is a correct next step to solve the equation? x – 4 = –7 x + 4 = –7 –x – 4 = 7 x + 4 = 7

OpenStudy (anonymous):

Add four to each side..

OpenStudy (anonymous):

So that the answer would be x=11

OpenStudy (anonymous):

howugetthat

OpenStudy (anonymous):

How did I get 11?

OpenStudy (anonymous):

|x-4| = 7 +4 +4 ^^Add four to each side, so that the "x" is alone on one side and the whole number on the other side.^^ Now you have... x= 7+4 Simplify it by adding both numbers... So x= 11

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Np